OpenStudy (anonymous):

Solve the inequality (x+4)/(x-5) <=0 Write your answer in interval notation Answer: -4<=x<5 Interval notation: [-4,5)

OpenStudy (anonymous):

Is this answer correct? @stamp I used your wolfram.com link you gave me, to double check it, but I want to be 100% by polling the audience. And to make sure the interval notation was correct.

OpenStudy (anonymous):

This is correct, yes! Although in Finland we use the interval notations [-4,5[. So I don't know about America, sorry! :)

OpenStudy (stamp):

How are you supposed to solve this one

OpenStudy (stamp):

I can see that (x+4) <= 0 leads to x <= 4

OpenStudy (anonymous):

I'm glad I'm not in Finland

OpenStudy (anonymous):

I don't know?

OpenStudy (stamp):

Ohhhh wait

OpenStudy (stamp):

If it is less than zero, it must be negative so x+5 <= 0 x <= -5 but x =/= 5 so x < 5

OpenStudy (stamp):

Why is it x >= 4

OpenStudy (anonymous):

if you graph it on a number line it looks like

OpenStudy (stamp):

well if it was lessthan you would get -/- which is +

OpenStudy (stamp):

but if its greater than you have +/- = - is <0

OpenStudy (anonymous):

You have to draw the functions (x+4 and x-5) and check when they're <= 0. :) I could do that, but I'm too lazy, haha :D

OpenStudy (anonymous):